The quality-control inspector of a production plant will reject a batch of syringes if two or more defective syringes are found in a random sample of eight syringes taken from the batch. Suppose the batch contains 5% defective syringes. A button hyperlink to the SALT program that reads: Use SALT.

(a) Make a histogram showing the probabilities of r = 0, 1, 2, 3, , 7 and 8 defective syringes in a random sample of eight syringes.

(b) Find μ. (Enter your answer to two decimal places.) μ = syringes What is the expected number of defective syringes the inspector will find? (Enter your answer to two decimal places.) syringes

(c) What is the probability that the batch will be accepted? (Round your answer to three decimal places.) (

d) Find σ. (Round your answer to three decimal places.) σ = syringes
Answer #1
here this is binomial with parameter n=8 and p=0.05

a)

P(x) 0.70000 0.60000 0.50000 0.40000 0.30000 0.20000 0.10000 0.00000 0 1 2 3 4 5 6 7 8 P(x)

b)

mean E(x)=μ=np=0.4

expected number of defective syringes the inspector will find =0.4

c)

probability that the batch will be accepted :

P(X>=2)=1-P(X<=1)= 1-∑x=0x-1   (nCx)px(q)(n-x) = 0.057

d)

standard deviation σ=√(np(1-p))=0.616